OpenStudy (anonymous):

solve the equation (x+4)(x-6)=21

OpenStudy (bahrom7893):

multiply it out and move everything to one side, then use the quadratic equation.

OpenStudy (imtant):

x^2-6x+4x-24=21 then bring 21 to the left, x^2-2x-24-21=0 x^2-2x-45=0 and factorize it,to get x..

OpenStudy (anonymous):

FOIL then bring 21 over then factor right?

OpenStudy (imtant):

ya..there will be two x values..